OpenStudy (anonymous):

find the derivative of (x-y)^2=y^2-xy

OpenStudy (anonymous):

something about implicit differentiation i think

OpenStudy (amistre64):

implicit rules are the same as explicit rules

OpenStudy (amistre64):

\[D[(x-y)^2=y^2-xy]\] \[D[(x-y)^2]=D[y^2]-D[xy]\] \[2(x-y)*D[x-y]=2yy'-(D[x]y+xD[y])\] \[2(x-y)*(D[x]-D[y])=2yy'-(x'y+xy')\] \[2(x-y)(x'-y')=2yy'-(x'y+xy')\]

OpenStudy (amistre64):

when you determine what this is "with respect to" then you can determine what to do with the ' parts

OpenStudy (anonymous):

it says to find dy/dx

OpenStudy (amistre64):

dy/dx = y' dx/dx = x' = 1

OpenStudy (amistre64):

\[2(x-y)(x'-y')=2yy'-(x'y+xy')\] \[2(x-y)(1-y')=2yy'-(y+xy')\] do the algebra to solve for y'

OpenStudy (amistre64):

the key tho, is to realize the implicits are no different than explicits; the rules dont change, just the manner in which your used to seeing them changes - basically because they teach this stuff in a method manner instead of a comprehesive manner

OpenStudy (anonymous):

i got the formula and the rules right, but when i solve for y' i keep getting the wrong answer

OpenStudy (amistre64):

yeah, algebra can be a pain :) can you show me your steps?

OpenStudy (anonymous):

(2x-2y)(1-y')=2yy'-y-xy' 2x-2xy'-2y+2yy'=2yy'-y-xy' xy'=y+2x-2y xy'=2x-y y'=(2x-y)/x oh nevermind i got it :)

OpenStudy (anonymous):

thank you so much

OpenStudy (amistre64):

yay!! :)